A man takes 16 breath per minute at 18degree Celsius and 760mmhg. What is his rate of his breathing on a mountain top with a temperature of -5 degrees and the pressure is 240mmhg. Assume that he needs the same mass of air per minute


1
Expert's answer
2021-01-30T16:48:09-0500
"\frac{p}{NT}=const\\N_2=N_1\frac{p_2T_1}{p_1T_2}
\\N_2=16\frac{240\cdot(18+273)}{760\cdot(-5+273)}=5.5"

A man needs 5.5 breath per minute.